What is the probability of selecting a person with the last two digits of their telephone number being 5 and 8? Hint: look at digits in this way _ _ _-_ _ _-_ _ 58. Is this with or without replacement? Explain.
Answer by stanbon(75887)

What is the probability of selecting a person with the last two digits of their telephone number being 5 and 8?
Hint: look at digits in this way _ _ _-_ _ _-_ _ 58. Is this with or without replacement?
-----
With Replacement:
Prob = 10^8/10^10 = 1/100
------------------------------
Without Replacement:
8!/10! = 1/(10*9) = 1/90
==================================
Cheers,
Stan H.
=============